two solid sphere of same material have radii in the ratio 1:2.Their moment of inertia in the ratio

Grade:12

1 Answers

Krishan Pareek
6 years ago
answer will be M*R*R/m*r*ras mass depends on volume and volume depends on cube of radiusso mass depends on r*r*rso ratio becomes (R/r)5= 1/32ratio of inertia = 1:32